OpenStudy (anonymous):

Which of the following is a factor of 32x^3 + 4?

OpenStudy (saranya):

4 + 32x^3 = 4*(1 + 8x^3) = 4*(1 + 2x)*(1 - 2x + 4x^2)

OpenStudy (saranya):

To be more clear :) 32x^3 + 4 =4(8x^3 + 1) =4((2x)^3 + 1^3) =4(2x + 1)((2x)^2 - 2x + 1) =4(2x +1)(4x^2 - 2x + 1)

OpenStudy (anonymous):

Cheers!